OpenStudy (anonymous):

Here's an interesting question for you all that almost tricked me at first. It's from a recent SAT exam. A hiking trail is 19 miles long. There is a marker at the start of the trail and a marker every 2.5 miles. There is a rest area at the start of the trail and a rest area every 6 miles. What is the ratio of the number of markers to the number of rest areas on the entire trail? (A) 2 to 1 (B) 4 to 1 (C) 5 to 12 (D) 7 to 3 (E) 8 to 3 Note that you should try to complete it in a realistic amount of time... on the actual exam one usually doesn't have much more than 5 (if you finish everything else rather quickly you might have maybe 9 or 10) minutes to dedicate to a particular math problem.

OpenStudy (anonymous):

1+19/2.5 is close to 10 . 1+19/6 is close to 4. closest answer to that is E. i remember this question last year. :/

OpenStudy (mathstudent55):



OpenStudy (ganpat):

for Marker: 19/ 2.5 FOr Rest area: 19/6 therfore, ratio of Number of markers to No of rest = 6/ 2.5.. i guess, it should be 2 : 1... roughly !!!

OpenStudy (anonymous):

-.- alright then. good thing i wasn't taking the sat at midnight. :/

OpenStudy (anonymous):

The answer boils down to: $$\frac{\lceil\frac{19}{2.5}\rceil}{\lceil\frac{19}{6}\rceil}=\frac84=\frac21$$

OpenStudy (anonymous):

@Ganpat you're neglecting both that one can't have a fraction of a marker/rest-area as well as that there are both a marker and rest-area at the beginning :-)

OpenStudy (ganpat):

yeah right.. so i rounded up in the end.. well, i know its not a good practise for large numbers...

OpenStudy (anonymous):